# Jose Tupayachi

### v1

@@ 5.1 Prediction Analysis
To evaluate the performance of each configuration [^2] https://github.com/XXXXX/blob/main/results.tex, we analyzed the average error metrics across combinations of lag hours, number of clusters, sequence lengths, and models. A visual summary of these relationships is provided in Figures 3 and 4.
@@
Shorter lag hours, particularly LAG = 1, tend to yield lower MAE and RMSE for most models, as clearly shown in Figure 3. For example, 1DCNN achieves its lowest MAE of 0.06 and RMSE of 0.52 at this lag. Increasing the number of clusters generally improves 1DCNN performance, whereas RNN and GRU models show more sensitivity to cluster size, with the optimal 1DCNN configuration corresponding to CLUSTERS = 10.
Figure 4: Comparison of TWGCN models against baseline machine learning models across error metrics (MAE, RMSE, MSE, SMAPE). TWGCN models are shown as boxplots, while baseline ML models are shown as overlaid stripplots for direct comparison.
Moderate sequence lengths, typically between 4 and 8, often produce lower SMAPE values, indicating better relative forecasting accuracy; for instance, 1DCNN attains a balanced performance with SEQ_LENGTH = 8. Across all models, 1DCNN frequently outperforms LSTM, RNN, and GRU in terms of RMSE and MAE, particularly for configurations with LAG = 1 and CLUSTERS = 10. Overall, the lowest MAE (0.06) and RMSE (0.52) are achieved by 1DCNN using the configuration LAG = 1, CLUSTERS = 10, and SEQ_LENGTH = 8, suggesting that this setup best balances spatial clustering and temporal information. Full results extract can be found under (Tables 3 and 4)  [^3] Note: L = LAG_HOURS, C = NUM_CLUSTERS, $\alpha$ = ALPHA, SL = SEQ_LENGTH.
Table 3: Top 5 configurations per model type
Model L C $\alpha$ SL MAE MSE RMSE SMAPE
1DCNN 3 5 0.66 8 0.064 0.328 0.528 5.9%
1 10 0.66 8 0.064 0.330 0.520 5.8%
3 5 0.33 8 0.068 0.336 0.535 6.4%
3 5 0.33 4 0.068 0.336 0.511 6.5%
6 10 0.5 8 0.069 0.355 0.555 6.3%
RNN 6 5 0.33 12 0.082 0.549 0.692 4.3%
3 10 0.66 8 0.083 0.504 0.663 3.6%
3 5 0.5 12 0.084 0.512 0.673 4.2%
1 5 0.66 12 0.086 0.510 0.673 4.2%
3 5 0.5 4 0.086 0.474 0.632 4.1%
LSTM 3 5 0.33 12 0.082 0.464 0.641 4.2%
6 5 0.5 8 0.088 0.508 0.661 5.0%
6 10 0.5 8 0.092 0.509 0.666 5.7%
1 5 0.66 4 0.093 0.470 0.629 5.2%
1 5 0.66 12 0.094 0.520 0.682 5.0%
GRU 3 5 0.33 12 0.102 0.544 0.693 6.0%
1 10 0.33 8 0.116 0.553 0.693 7.1%
1 10 0.66 12 0.117 0.583 0.721 8.1%
1 5 0.66 8 0.121 0.605 0.727 7.4%
3 10 0.33 8 0.124 0.628 0.746 8.0%
Table 4: Baseline models performance
Model MAE MSE RMSE SMAPE
@@
The analysis in Figures 3 and 4 highlights a clear separation between classical baselines and deep temporal–graph architectures. The boxplots illustrate that TWGCN-based models not only achieve consistently lower errors across all metrics but also display reduced variability compared to baseline methods. This robustness is especially relevant for practical deployment, where stable performance across different configurations is as critical as peak accuracy. In contrast, baseline machine learning models (Table 4) exhibit both higher errors and wider dispersion, reflecting their limited ability to capture nonlinear temporal and spatial dependencies inherent in the data.
Equally important is the comparative behavior among the deep learning architectures themselves, as summarized in Table 3. While convolutional structures capturing localized temporal dependencies, recurrent models demonstrate more nuanced sensitivity to hyperparameters such as lag depth and sequence length. The distributional plots in Figure 3 further emphasize how these differences manifest across error scales. Together, these results suggest that the performance gap is not simply a matter of choosing the proposed TW-GCN over classical methods, but of selecting architectures that best exploit temporal and spatial dynamics in tandem.
6 Infrastructural Challenges
Based on the results presented in section 5 we explore three optimized model and configurations to assess the impact of different temporal horizons and clustering strategies on the predictive performance of EV charging station usage in Tennessee First 1-hour lag time (LAG_HOURS = 1) and divides the data into 10 clusters (NUM_CLUSTERS = 10). It employs a 1DCNN with a sequence length (SEQ_LENGTH) of 4 and an alpha ($\alpha$) value of 0.66, which balances different loss components or regularization terms.
While a 1-hour lag can capture immediate fluctuations, it may be overly sensitive to noise and transient anomalies, making it less reliable for modeling sustained behavioral trends. In contrast, 3- and 6-hour lags smooth out short-term volatility, better reflecting underlying patterns and enabling more stable predictions, critical for operational planning (e.g., shift scheduling, energy load balancing) and strategic decisions (e.g., infrastructure adjustments, demand forecasting).
Three hours lags increases the lag time to 3 hours while reducing the number of clusters to 5, maintaining the same 1DCNN architecture, sequence length of 4, and alpha value of 0.66.
Another lag configuration that relies on 6 hours and reverts to 10 clusters, again using the 1DCNN with a sequence length of 4 and $\alpha$ = 0.66. Across all three models, the 1DCNN and sequence length remain consistent, suggesting a preference for this architecture in capturing temporal patterns, while variations in lag hours and cluster numbers explore different temporal and data partitioning strategies. The consistent alpha value indicates a stable weighting mechanism across configurations.
@@ 6.1 Implications for Decision Makers
Our study offers valuable insights for policymakers, urban planners, utility providers, and investors engaged in EV infrastructure planning. By utilizing spatial-temporal graph neural networks integrated with long short-term memory models, we enable accurate medium-term predictions of EV charging demand based on spatial, weather-related, and traffic variables. These predictive capabilities facilitate optimized infrastructure deployment, helping to prevent over- or under-investment in charging stations, particularly in slow-adoptant regions.
Furthermore, the findings support more resilient power grid management by forecasting load distributions and peak usage times. For investors, this modeling improves decision-making efficiency by identifying high-potential areas for deployment. The study also contributes to broader sustainability efforts by addressing the uneven spatial distribution of EV infrastructure, promoting equitable access, and supporting the reduction of transportation-related greenhouse gas emissions. Therefore, our approach can guide the creation of adaptive policies and incentive structures that are responsive to evolving patterns of demand and adoption.
Subsequently, we examine how traffic conditions, weather, and POIs influence EV charging patterns, analyzing their relationships across the three main sources of information.
@@ 6.1.1 Traffic
By uniting the station dataset with the TMC traffic data, we were able to contextualize charging demand within the surrounding network. The nearest-neighbor matching approach allowed each station observation to be associated with the traffic conditions of its closest road segments at the same point in time. This fusion revealed not only how traffic speed and travel time vary around charging locations, but also how reliability indicators such as delay per mile, confidence score, and reference speed influence accessibility. Analyzing these combined features provides key insights into the interaction between road congestion and charging station utilization, to detect the importance of traffic-aware planning in understanding EV demand patterns. We compute the delay for each road segment relative to the reference speed . The analysis of traffic delays and EV charging behavior for the month of January (Figure 5) reveals that the overall correlation between delays and charging energy is weak indicating that traffic congestion alone does not strongly influence charging demand.
Figure 5: Traffic delays at EV stations based on neighboring TMCs (January 2024). The figure presents the average delay (km/h) for each EV station, computed from the 1, 5, and 10 nearest TMC segments. Delay is defined as the difference between the reference (free-flow) speed and the observed TMC speed. Each dot represents the mean delay for the corresponding spatial aggregation level, highlighting how traffic conditions in the vicinity of EV stations vary with proximity. Certain stations exhibit greater sensitivity to surrounding traffic, showing higher variability in delay across neighboring TMCs.
Figure 6: Spatial comparison of EV charging energy consumption (top) and traffic delays (bottom) across Tennessee, with density shading and station-level values overlaid.
However, a statistically significant difference in charging patterns is observed across delay segments. Low-delay stations record higher average energy consumption (6.45 kWh), while high-delay stations show substantially lower averages (4.76 kWh), with a mean difference of 1.69 kWh ($p<0.001$). This suggests that in congested areas, drivers tend to engage in shorter, opportunistic charging sessions rather than longer, high-capacity charges. From an executive perspective, infrastructure investments should prioritize larger-capacity charging hubs in low-delay zones, where demand is stronger and more consistent, while in high-delay zones, strategies should emphasize fast chargers designed for shorter sessions and higher turnover. While traffic delays are not a direct predictor of charging demand, they affect driver behavior and charging patterns. In our framework, this information forms part of the input to the stacked model used in the TW-GCN, enabling the network to better capture the interplay between traffic conditions and charging behavior in January and to guide resource allocation more efficiently.
@@
Spatial analysis revealed stations can form up to 29 geographic clusters, with only 2 stations being isolated. This clustering highlights the concentration of charging infrastructure in high-demand areas, likely reflecting population density and urban planning considerations. When categorizing stations based on performance, 468 events of 2,520 were identified as high-energy, low-delay (optimal), whereas 317 events were high-delay, low-energy (inefficient). Optimal charging behavior are typically situated in locations that facilitate quick and frequent charging, while inefficient charging operations may face accessibility issues or lower utilization.
Regional differences (Figure 6) were also observed: the northern region, with 1237 charging events, exhibited higher average energy consumption (7.56 kWh) and slightly higher delays (2.07 km/h), whereas the southern region, with 1285 stations, showed lower energy consumption (4.69 kWh) and shorter delays (1.24 km/h). This suggests that northern stations experience higher local demand and potentially longer dwell times. Analysis of station proximity further revealed that among station pairs separated by less than 5.5 km, the average energy difference was only 3.80 kWh, indicating consistent usage patterns among nearby stations and supporting the potential for coordinated load management in clustered areas.
These findings indicate that energy consumption at EV charging stations is only weakly related to traffic delays, but strongly influenced by spatial clustering and regional demand patterns. Stations performing optimally tend to be located in high-demand areas with minimal congestion, and nearby stations exhibit similar energy usage, highlighting opportunities for targeted infrastructure planning and efficiency improvements.
@@ 6.1.2 Weather
January, being a winter month, features variable conditions with temperatures ranging from -1°F to 78°F, humidity reaching up to 100%, and wind speeds up to 32 mph. Correlation analysis indicates that energy demand is only weakly associated with weather variables, suggesting that EV usage is largely driven by human behavior rather than environmental conditions. However, daily weather patterns can still subtly influence charging: humid or rainy days may reduce trip frequency or encourage shorter charging sessions, while dry and mild days may promote longer trips and slightly higher energy consumption. Extremely cold days could slightly increase energy usage due to cabin heating or battery preconditioning. Overall, while weather introduces minor fluctuations, the primary determinants of energy consumption are user schedules, commuting patterns, and station availability throughout January.
@@ 6.1.3 Points of Interest
Energy consumption was analyzed across various POIs, including hospitals, supermarkets, schools, restaurants, and parks. Overall, hospitals, schools, and parks exhibit similar statistics, with a mean energy consumption of approximately 6.10 kWh, a median of 1.84 kWh, and a standard deviation of 10.33 kWh across 2,522 observations, totaling around 15,381 kWh. Supermarkets show a more heterogeneous pattern: most charging operation (2,519) have a mean energy of 6.09 kWh, while a small subset of three locations exhibits higher consumption (9.68 kWh), suggesting potential high-consumption outliers. Restaurants also display variability, with the majority (2,498 locations) averaging 6.14 kWh, but a smaller group of 24 locations showing significantly lower usage (1.92 kWh). Statistical testing indicates that the presence of supermarkets is associated with higher mean energy consumption compared to locations without them, although this difference is not statistically significant ($t=0.74$, $p=0.54$). In contrast, locations with restaurants consume significantly less energy than those without (t=10.57t=-10.57, $p<0.001$). An interaction analysis between hospitals and supermarkets reveals that locations without both POIs (n = 2,519) have a mean energy of 6.09 kWh, while locations without hospitals but with supermarkets (n = 3) have a higher mean of 9.68 kWh, indicating a potential combined effect, though limited by small sample size. Overall, these results highlight that energy consumption patterns vary across POIs, with restaurants and supermarkets showing the most notable deviations.
@@ 6.2 Prediction Horizon Analysis
This section examines how prediction horizon and regional characteristics affect model performance. By analyzing RMSE across Tennessee’s ZIP codes, we explore the trade-offs between responsiveness and stability in forecasts. Figure 8 illustrates the geographic distribution of RMSE using a 3-hour lag model, which was identified as the best-performing configuration in our analysis. Key parameters for this model are: $\texttt{LAG}=3\penalty 10000\ \text{hours}$, $\texttt{SEQ\_LENGTH}=4$, and $\texttt{CLUSTERS}=10$. Notably, the station count exhibits almost no correlation with RMSE ($r=0.008$), indicating that simply increasing the number of stations does not guarantee better predictive accuracy. In addition to benchmark metrics, we also report several other metrics to provide a more comprehensive evaluation of the model’s performance. The 1D-CNN model achieves the following:
  • Coefficient of Determination ($R^{2}$): 0.9659
  • Mean Absolute Scaled Error (MASE): 0.0238
  • Explained Variance Score: 0.9661
Figure 8: Geographic distribution of prediction errors (RMSE) across Tennessee’s Grand Divisions using a 3-hour lag model. ZIP codes are colored by RMSE, with regional averages and standard deviations annotated. Major cities are marked for reference, showing how model performance varies spatially across East, Middle, and West Tennessee.
To better understand model performance and infrastructure needs, we examine regional insights across the state of Tennessee:
  • East Tennessee Encompasses 22 ZIP codes with a total population of 404,127. It has moderate RMSE ($0.486\pm 0.473$) but low station density (0.82 per 10,000 people). High-population ZIPs such as 37876 (28,66128{,}661 residents) exhibit elevated RMSE (0.769), highlighting potential infrastructure gaps. Conversely, ZIPs 37620 and 37814 show low RMSE (0.36 and 0.03, respectively), indicating strong model fit and suitability for targeted deployment or expansion of services.
  • Middle Tennessee includes 19 ZIP codes and roughly 294,000 residents. Despite the highest station density (2.31 per 10,000), it has the highest average RMSE (0.497). Specific ZIPs, including Carthage and McMinnville, record RMSE above 1.0, suggesting irregular usage patterns or demand fluctuations that challenge predictive accuracy.
  • West Tennessee Covers 8 ZIP codes with a population of 130,091. It maintains moderate RMSE (0.487) and station density (1.92 per 10,000), yet variability remains high (SD = 0.516). ZIPs such as 38340 and 38320 have RMSE exceeding 0.74, reflecting local and regional behaviors.
  • Mixed undefined ZIPs (4 entries) exhibit the lowest average RMSE (0.463) and the highest station density (2.40 per 10,000). Limited entries and undefined geography may provide useful reference points.
Short prediction horizons (1-hour) produce highly responsive forecasts suitable for near-term operational decisions. They allow the model to quickly adapt to recent patterns and anomalies but can introduce volatility due to limited long-term visibility. Extending the horizon to 3 hours yields the best trade-off between responsiveness and stability, producing low RMSE across most regions. Longer horizons (6 hours) slightly reduce responsiveness and may increase errors in Middle and West Tennessee, while East Tennessee maintains low RMSE (0.426), demonstrating scalable predictive accuracy. Combining RMSE, population, and station density data supports evidence-based decision making. Stakeholders can prioritize infrastructure in high-population, low-RMSE areas, investigate high-error regions for demand anomalies, and refine models to balance responsiveness, accuracy, and equitable deployment across Tennessee’s EV ecosystem.

### v3

@@ 5.1 Prediction Analysis
To evaluate the performance of each configuration [^1] https://github.com/XXXXXXXXX/XXXXXXXXX/blob/main/Readme.md (MASKED FOR PEER REVIEW), we analyzed the average error metrics across combinations of lag hours, number of clusters, sequence lengths, and models. A visual summary of these relationships is provided in Figures 2 and 3.
@@
Figure 3: Comparison of TW-GCN models against baseline machine learning models across error metrics (MAE, RMSE, MSE, SMAPE). TW-GCN models are shown as boxplots, while baseline ML models are shown as overlaid stripplots for direct comparison.
Shorter lag hours, particularly LAG = 1, tend to yield lower MAE and RMSE for most models, as clearly shown in Figure 2. For example, 1DCNN achieves its lowest MAE of 0.064 and RMSE of 0.575 at this lag. Increasing the number of clusters generally improves 1DCNN performance, whereas RNN and GRU models show more sensitivity to cluster size, with the optimal 1DCNN configuration corresponding to CLUSTERS = 10.
Moderate sequence lengths, typically between 4 and 8, often produce lower SMAPE values, indicating better relative forecasting accuracy; for instance, 1DCNN attains optimal performance with SEQ_LENGTH = 8, suggesting that capturing recent temporal patterns without excessive historical context improves predictive reliability. Across all models, 1DCNN frequently outperforms LSTM, RNN, and GRU in terms of RMSE and MAE. Overall, the lowest MAE (0.064) and RMSE (0.573) are achieved by 1DCNN using the configuration LAG = 3, CLUSTERS = 5, and SEQ_LENGTH = 8, suggesting that this setup best balances spatial clustering and temporal information. Full results extract can be found under (Tables 4 and 5)  [^2] Note: L = LAG_HOURS, C = NUM_CLUSTERS, $\alpha$ = ALPHA, SL = SEQ_LENGTH.
Table 4: Top 5 configurations per model type. (↓ indicates lower is better)
Model L C $\alpha$ SL MAE ↓ MSE ↓ RMSE ↓ SMAPE ↓
1DCNN 1 10 0.66 8 0.064 0.330 0.575 0.058
3 5 0.66 8 0.064 0.328 0.573 0.059
3 5 0.33 4 0.068 0.336 0.579 0.065
3 5 0.33 8 0.068 0.336 0.579 0.064
6 10 0.5 8 0.069 0.355 0.596 0.063
RNN 6 5 0.33 12 0.082 0.549 0.741 0.043
3 10 0.66 8 0.083 0.504 0.710 0.036
3 5 0.5 12 0.084 0.512 0.715 0.042
1 5 0.66 12 0.086 0.510 0.714 0.042
3 5 0.5 4 0.086 0.474 0.689 0.041
LSTM 3 5 0.33 12 0.082 0.464 0.681 0.042
6 5 0.5 8 0.088 0.508 0.713 0.050
6 10 0.5 8 0.092 0.509 0.714 0.057
1 5 0.66 4 0.093 0.470 0.685 0.052
1 5 0.66 12 0.094 0.520 0.721 0.050
GRU 3 5 0.33 12 0.102 0.544 0.738 0.060
1 10 0.33 8 0.116 0.553 0.743 0.071
1 10 0.66 12 0.117 0.583 0.763 0.081
1 5 0.66 8 0.121 0.605 0.778 0.074
3 10 0.33 8 0.124 0.628 0.793 0.080
Table 5: Baseline models performance. (↓ indicates lower is better)
Model MAE ↓ MSE ↓ RMSE ↓ SMAPE ↓
@@
The analysis in Figures 2 and 3 highlights a clear separation between classical baselines and deep temporal–graph architectures. The boxplots illustrate that TW-GCN based models not only achieve consistently lower errors across all metrics but also display reduced variability compared to baseline methods. This robustness is especially relevant for practical deployment, where stable performance across different configurations is as critical as peak accuracy. In contrast, baseline machine learning models (Table 5) exhibit both higher errors and wider dispersion, reflecting their limited ability to capture nonlinear temporal and spatial dependencies inherent in the data.
Equally important is the comparative behavior among the deep learning architectures themselves, as summarized in Table 4. While convolutional structures capture localized temporal dependencies, recurrent models demonstrate more nuanced sensitivity to hyperparameters such as lag depth and sequence length. Figure 2 further emphasize how these differences manifest across error scales. Together, these results suggest that the performance gap is not simply a matter of choosing the proposed TW-GCN over classical methods, but of selecting architectures that best exploit temporal and spatial dynamics.
6 Analysis & Insights
Based on the results presented in Section 5, we explore three optimized models and configurations to assess the impact of different temporal horizons (1, 3, 6) hours and clustering strategies on the predictive performance of EV charging station usage in Tennessee. For example, a 1-hour lag time (LAG_HOURS = 1) combined with 10 clusters (NUM_CLUSTERS = 10), a sequence length (SEQ_LENGTH) of 4, and an alpha ($\alpha$) value of 0.66 can effectively capture immediate fluctuations. However, this configuration may also be overly sensitive to noise and transient anomalies, making it less reliable for modeling sustained behavioral trends. In contrast, 3- and 6-hour lags smooth out short-term volatility, better reflecting underlying patterns and enabling more stable predictions, critical for operational planning (e.g., shift scheduling, energy load balancing) and strategic decisions (e.g., infrastructure adjustments, demand forecasting).
A three hour lag increases the delay to 3 hours while reducing the number of clusters to 5, maintaining the same 1DCNN architecture, a sequence length of 4, and an $\alpha$ value of 0.66. Another configuration uses a 6-hour lag and reverts to 10 clusters, again employing the 1DCNN with a sequence length of 4 and $\alpha=0.66$. Across all three models, the 1DCNN and sequence length remain consistent, suggesting a preference for this architecture in capturing temporal patterns, while variations in lag hours and cluster numbers explore different temporal and data partitioning strategies. The consistent alpha value indicates a stable weighting mechanism across configurations.
@@ 6.1 Implications for Decision Makers
Our study offers valuable insights for policymakers, urban planners, utility providers, and investors engaged in EV infrastructure planning. By utilizing TW-GCN, we enable accurate medium-term predictions of EV charging demand based on spatial, weather-related, and traffic variables. These predictive capabilities facilitate optimized infrastructure deployment, helping to prevent over or under investment in charging stations, particularly in slow adopting regions.
Furthermore, the findings support more resilient power grid management by forecasting load distributions and peak usage times. For investors, this modeling improves decision making efficiency by identifying high-potential areas for deployment. The study also contributes to broader sustainability efforts by addressing the uneven spatial distribution of EV infrastructure, promoting equitable access, and supporting the reduction of transportation greenhouse gas emissions. Therefore, our approach can guide the creation of adaptive policies and incentive structures that are responsive to evolving patterns of demand and adoption. Subsequently, we examine how traffic conditions, weather, and POIs influence EV charging patterns, analyzing their relationships across the three main sources of information.
@@ 6.1.1 Traffic
By uniting the station dataset with the TMC traffic data, we were able to contextualize charging demand within the surrounding network. The nearest-neighbor matching approach allows each station observation to be associated with the traffic conditions of its closest road segments at the same point in time. This fusion revealed not only how traffic speed and travel time vary around charging locations, but also how reliability indicators such as delay per mile, confidence score, and reference speed influence accessibility. Analyzing these combined features provides key insights into the interaction between road congestion and charging station utilization to detect the importance of traffic-aware planning in understanding EV demand patterns. We compute the delay for each road segment relative to the reference speed. The analysis of traffic delays and EV charging behavior for January reveals that the overall correlation between delays and charging energy is weak, indicating that traffic congestion alone does not strongly influence charging demand.
However, a statistically significant difference in charging patterns is observed across delay segments. Low-delay stations record higher average energy consumption (6.45 kWh), while high-delay stations show substantially lower averages (4.76 kWh), with a mean difference of 1.69 kWh ($p<0.001$). This suggests that in congested areas, drivers tend to engage in shorter, opportunistic charging sessions rather than longer, high-capacity charges. From an executive perspective, infrastructure investments could prioritize larger-capacity charging hubs in low-delay zones, where demand is stronger and more consistent, while in high-delay zones, strategies should emphasize fast chargers designed for shorter sessions and higher turnover. While traffic delays are not a direct predictor of charging demand, they affect driver behavior and charging patterns. In our framework, this information forms part of the input to the stacked model used in the TW-GCN, enabling the network to better capture the interplay between traffic conditions and charging behavior in January and to guide resource allocation more efficiently.
@@
Spatial analysis revealed stations can form up to 29 geographic clusters, with only two stations being isolated. This clustering highlights the concentration of charging infrastructure in high-demand areas, likely reflecting population density and urban planning considerations. When categorizing stations based on performance, 468 events of 2,520 were identified as high-energy, low-delay (optimal), whereas 317 events were high-energy, low-energy (inefficient). Optimal charging behavior is typically situated in locations that facilitate quick and frequent charging, while inefficient charging operations may face accessibility issues or lower utilization.
Regional differences are also observed: the northern region, with 1,237 charging events, exhibited higher average energy consumption (7.56 kWh) and slightly higher delays (2.07 km/h), whereas the southern region, with 1285 stations, showed lower energy consumption (4.69 kWh) and shorter delays (1.24 km/h). This suggests that northern stations experience higher local demand and potentially longer dwell times. Analysis of station proximity further revealed that among station pairs separated by less than 5.5 km, the average energy difference was only 3.80 kWh, indicating consistent usage patterns among nearby stations and supporting the potential for coordinated load management in clustered areas.
These findings indicate that energy consumption at EV charging stations is only weakly related to traffic delays but strongly influenced by spatial clustering and regional demand patterns. Stations performing optimally tend to be located in high-demand areas with minimal congestion, and nearby stations exhibit similar energy usage, highlighting opportunities for targeted infrastructure planning and efficiency improvements.
@@ 6.1.2 Weather
January, being a winter month, features variable conditions with temperatures ranging from -1°F to 78°F, humidity reaching up to 100%, and wind speeds up to 32 mph. Correlation analysis indicates that energy demand is only weakly associated with weather variables, suggesting that EV usage is largely driven by human behavior rather than environmental conditions. However, daily weather patterns can still subtly influence charging: humid or rainy days may reduce trip frequency or encourage shorter charging sessions, while dry and mild days may promote longer trips and slightly higher energy consumption. Extremely cold days could slightly increase energy usage due to cabin heating or battery preconditioning. Overall, while weather introduces minor fluctuations, the primary determinants of energy consumption are user schedules, commuting patterns, and station availability throughout the studied period.
@@ 6.1.3 Points of Interest
Energy consumption was analyzed across various POIs, including hospitals, supermarkets, schools, restaurants, and parks. Overall, hospitals, schools, and parks exhibit similar statistics, with a mean energy consumption of approximately 6.10 kWh, a median of 1.84 kWh, and a standard deviation of 10.33 kWh across 2,522 observations, totaling around 15,381 kWh. Supermarkets show a more heterogeneous pattern: most charging operations (2,519) have a mean energy of 6.09 kWh, while a small subset of three locations exhibits higher consumption (9.68 kWh), suggesting potential high-consumption outliers. Restaurants also display variability, with the majority (2,498 locations) averaging 6.14 kWh, but a smaller group of 24 locations showing significantly lower usage (1.92 kWh). Statistical testing indicates that the presence of supermarkets is associated with higher mean energy consumption compared to locations without them, although this difference is not statistically significant ($t=0.74$, $p=0.54$). In contrast, locations with restaurants consume significantly less energy than those without (t=10.57t=-10.57, $p<0.001$). An interaction analysis between hospitals and supermarkets reveals that locations without both POIs (n = 2,519) have a mean energy of 6.09 kWh, while locations without hospitals but with supermarkets (n = 3) have a higher mean of 9.68 kWh, indicating a potential combined effect, though limited by small sample size. Overall, these results highlight that energy consumption patterns vary across POIs, with restaurants and supermarkets showing the most notable deviations.
@@ 6.2 Prediction Horizon Analysis
This section examines how the prediction horizon and regional characteristics affect model performance. By analyzing RMSE across Tennessee’s ZIP codes, we explore the trade-offs between responsiveness and stability in forecasts. Figure 5 illustrates the geographic distribution of RMSE using a 3-hour lag model, which was identified as the best-performing configuration in our analysis. Key parameters for this model are: $\texttt{LAG}=3\penalty 10000\ \text{hours}$, $\texttt{SEQ\_LENGTH}=4$, and $\texttt{CLUSTERS}=10$. Our analysis of EV charging stations across Tennessee (see Figures 5 and 6) can guide new site deployments. By analyzing station-level and regional forecast performance metrics, including RMSE, $R^{2}$, and sample size, we identify high-confidence deployment regions and recommend a phased rollout strategy to maximize return on investment and minimize operational risk. We provide a strategic analysis of EV charging stations across Tennessee to guide a partnership with ChargePoint for new site deployments. The analysis focuses on identifying both high-performing and underperforming stations and regions, highlighting opportunities and risks for executive decision-making.
Figure 5: Focused analysis of East Tennessee with best-performing station green star. Average stations are shown in yellow. Principal cities are labeled with blue triangles.
Figure 6: Performance analysis of individual charging stations using a 3-hour lag model and statewide view showing all stations colored by $R^{2}$ values.
Figure 7: Relationship between station usage (number of samples) and forecasting performance metrics (RMSE, $R^{2}$, and MAE) across Tennessee EV charging sites. Each subplot illustrates how predictive accuracy varies with data availability. Lower RMSE and higher $R^{2}$ values are observed at high-usage stations, indicating improved model reliability as sample size increases.
6.2.1 Regional Performance Insights
East Tennessee demonstrates the most reliable forecast performance, with the lowest RMSE among all grand divisions. This indicates minimal absolute error in predicted station usage, making it the most suitable region for immediate deployment. West Tennessee, while showing the highest $R^{2}$, also exhibits the highest RMSE, reflecting substantial uncertainty in forecasted usage. Deployments in this region carry higher operational and financial risk, and additional data collection is recommended before committing to large-scale investment. Middle Tennessee displays moderate RMSE and strong $R^{2}$, suggesting that it represents a promising but cautious expansion opportunity following successful pilot deployments in East Tennessee.
6.2.2 Station-Level Performance Insights
Several stations stand out for exceptional performance. The BrightRidge site in Johnson City achieves near-perfect prediction metrics, with an RMSE of $0.022$ and an $R^{2}$ of $0.999$, indicating minimal operational uncertainty and high confidence in forecasted demand. Similarly, the Jackson site exhibits extremely accurate forecasts, ranking among the top performers with consistently low RMSE and high $R^{2}$ values, making it ideal for early deployment and partnership initiatives. In contrast, a few stations display weaker predictive reliability. The Volunteer Parkway site in Bristol and a secondary BrightRidge location in Johnson City both present elevated RMSE and low $R^{2}$ values, suggesting substantial deviation between predicted and observed usage. These locations should be considered high-risk investments at present and warrant further observation or controlled pilot testing prior to full-scale deployment. From a system-wide perspective, correlation analysis reinforces these insights. The relationship between RMSE and sample count is mildly negative (r=0.175r=-0.175), consistent with the expectation that higher data availability improves model accuracy. Conversely, the correlation between $R^{2}$ and sample count is slightly positive ($r=0.078$), indicating that stations with more historical data tend to exhibit more reliable forecasts. Low-usage stations, which provide fewer observations, exhibit on average a $28.9\%$ higher RMSE compared to high-usage stations (see Figure 7). This highlights the importance of sustained data collection at low-traffic sites to stabilize predictions over time. All stations passed quality screening, confirming that model variance stems primarily from behavioral rather than data-integrity factors. Strategically, these results imply that high-usage, high-accuracy stations should serve as reference benchmarks for operational reliability, pricing calibration, and partnership modeling. Meanwhile, low-usage or high-variance sites should remain under watch or be integrated into limited pilot programs to validate demand potential before scaling investments. These findings, illustrated in Figure 7, provide a data-driven foundation for optimizing deployment sequencing, resource allocation, and collaborative expansion planning within Tennessee’s charging network.
6.2.3 Strategic Deployment Recommendations
Phase 1: Pilot Deployment
This phase should focus on the high-performing stations located in East Tennessee, where forecast reliability is the strongest. These sites present minimal operational risk and are ideal for testing early partnership models.
Phase 2: Secondary Expansion
The second phase should target stations in Middle Tennessee. Here, modular infrastructure can be deployed and scaled up gradually based on observed usage patterns and operational feedback.
Phase 3: Risk Assessment and Investigation
The final phase should include stations in West Tennessee and other low-performing sites. These areas require deeper analysis and targeted pilot projects to evaluate feasibility and mitigate risks before committing to full-scale investments.
6.2.4 Operational and Business-Level Actions
To optimize financial performance and operational efficiency, several strategic actions are recommended. Initial investments should prioritize high-performing stations characterized by low RMSE values and strong $R^{2}$ scores, ensuring reliable returns with minimal uncertainty. For moderate-performing sites, a modular deployment approach is advised to balance financial exposure and operational flexibility. Continuous monitoring of pilot sites will be essential to validate model forecasts and guide subsequent expansion decisions. Additionally, forming co-marketing or revenue-sharing partnerships at top-performing locations can enhance adoption and profitability. Investment in underperforming stations should be deferred until further historical data or improved predictive reliability becomes available.